OpenStudy (anonymous):

If a rectangle has a perimeter of 70, a length of x and a width of x - 9, find the value of the length of the rectangle. 4 31 13 22

OpenStudy (raden):

2(x + x-9) = 70 solve for x

OpenStudy (anonymous):

but there is no 70 @RadEn

OpenStudy (raden):

the perimeter known is 70, @liana123??

OpenStudy (anonymous):

ik but what one would it be 4 31 13 22

OpenStudy (raden):

2(x + x-9) = 70 can you solve for x in the equation above ?

OpenStudy (anonymous):

me

OpenStudy (anonymous):

if u talking about her maybe she can't cause she asked my question @RadEn

OpenStudy (anonymous):

shes saying me and thx

OpenStudy (raden):

ahh, i saw the same pic's profile, but your name is different. haha :p

OpenStudy (anonymous):

I dont have a profile pic @RadEn

OpenStudy (anonymous):

70 divide -18 ???

OpenStudy (raden):

i assumed that's an owl pic :) 2(x + x-9) = 70 divided by 2 on both sides, get x + x-9 = 35 2x - 9 = 35 2x = 35 + 9 2x = 44 x = 44/2 = 22

OpenStudy (anonymous):

thx for help

OpenStudy (anonymous):

anytime lol

OpenStudy (anonymous):

not you